OpenStudy (anonymous):

Find the derivative of f(x) = 4x + 7 at x = 5

OpenStudy (agent0smith):

Can you find the derivative of f(x) first?

OpenStudy (agent0smith):

This is from a precalc class right?

OpenStudy (anonymous):

@agent0smith yes c:

OpenStudy (anonymous):

im confused

OpenStudy (agent0smith):

Use f'(x) = (f(x+h) - f(x))/h familiar with this...? Equation editor is not working btw

OpenStudy (anonymous):

sorta

OpenStudy (agent0smith):

first, f(x+h) = 4(x+h) +7 then use the above formula f'(x) = ( 4(x+h) + 7 - (4x+7) )/h

OpenStudy (anonymous):

do i start solving?

OpenStudy (agent0smith):

Yep, simplify

OpenStudy (anonymous):

so 4?

OpenStudy (agent0smith):

Yep!

OpenStudy (e.mccormick):

Just as a thing to remember, the derivative is the slope of the tangent line. Well, if something is a straight line, it is its own tangent line. If it is curved there is more to it, but for simple equations with no powers or rational equations involved, you are looking for the slope.

OpenStudy (agent0smith):

^yeah, but with the wording of the question, i think they're looking for use of that method.

OpenStudy (e.mccormick):

You know better than me, seeing as it is what you do! I'm just adding in some food for thought.

OpenStudy (agent0smith):

Yeah, and you know you've done the work right if you get the slope of the line.